OpenStudy (waheguru):

1. A 40.0 kg child sits on a swing 50.0 cm above the ground. Her mother pushes her to a height of 1.2 m above the ground. What is the maximum speed of the girl on the swing?

OpenStudy (loser66):

to me, potential energy = kinetic energy (1/2)mgh = (1/2) mv^2 v^2 = gh and h = 1.2m - 0.5m = 1.15 m take square root to get v

OpenStudy (waheguru):

why does height = 1.2 - 0.5

OpenStudy (waheguru):

I was using the longer formula where Kinetic Energy Initial + Potential Energly Initial = Kin Energy FINal + Pot Energy Final

OpenStudy (loser66):

because originally, the child is at 0.5m above the ground, not on the ground

OpenStudy (loser66):

the same,

OpenStudy (waheguru):

no that way I get a negative number and I cant root it

OpenStudy (waheguru):

I get root of -13.72 but that is not possible

OpenStudy (loser66):

before moving, the potential energy =0 (didn't move yet) after moving, the kinetic energy =0

OpenStudy (loser66):

the negative sign show that it is upward, no meaning for the value.

OpenStudy (waheguru):

are you sure?

OpenStudy (loser66):

I think so.

OpenStudy (waheguru):

can you explain @amistre64

OpenStudy (waheguru):

Why is there a negative value?

OpenStudy (amistre64):

the question is ambiguous

OpenStudy (waheguru):

what do you mean?

OpenStudy (amistre64):

i mean, what time intervals are involved? when do we start calculating the speed? is the mothers pushing with all her strength or is this just getting her up to a certain height and letting go?

OpenStudy (waheguru):

The question is all that is given, btw this is grade 11 physics

OpenStudy (amistre64):

then we can prolly assume that the girl is pushed to a given height and let go an dallowed to jsut fall along the arced path.

OpenStudy (waheguru):

Its sort of like a pendulum

OpenStudy (amistre64):

at the bottom of the swings arc, all of the stored energy from being lifted to a given height is used, and we are in full kinetic mode

OpenStudy (amistre64):

assuming no loss of energy: K = P at the bottom of the arc, so it must be equal to the full potential to start with, and K = 1/2 mv^2 as pointed out